OpenStudy (anonymous):

I need someone good at math Please tell me how to factor to solve: (2x-3)(3x+2)≤ 0. The answer to this equation is (-2/3) ≤ x ≤ (3/2)

OpenStudy (anonymous):

okay but the other side is whats tricky

OpenStudy (anonymous):

how can u you just randomly split the multiplication problem into two equations like thiat?

OpenStudy (anonymous):

plus thats wrong and does not match me answer!!! wow no one know show to do this1

OpenStudy (anonymous):

\[(2x-3)(3x+2) \leq 0\] \[(2x-3)= 0\] \[(2x)= 3\] \[(x)= 3/2\] \[(3x+2)= 0\] \[(3x)= -2\] \[(x)= -2/3\] We have to check a points less than -2/3, a points between -2/3 and 3/2 and finally a points bigger than 3/2

OpenStudy (anonymous):

Randomly check x= -1 0 2